# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

### v1

@@ A.4 Further results
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 8.4 0.6 2.8 9.7 0.5 0.4 11.1 0.4 0.6 10.5 0.6 0.6 10.1 0.5 1.3 12.0 0.4 0.5 11.9 0.9 3.6 12.0 0.9 3.8 13.5 0.9 2.7 12.7 0.8 4.9 11.3 0.9 1.2 14.2 0.6 3.6
BeerRNN 11.7 0.8 4.7 8.0 0.5 0.5 3.3 0.2 0.4 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.4 12.5 0.9 2.2 12.1 0.9 1.8 12.6 0.8 1.4 13.0 0.9 1.9 14.1 0.6 4.1
Bouali 4.1 0.1 0.2 8.3 0.5 0.0 7.3 0.3 0.1 8.6 0.1 0.0 11.1 0.3 0.1 13.2 0.5 0.1 10.6 0.8 0.9 9.6 0.7 1.1 12.2 0.7 1.2 12.3 0.8 0.8 12.9 0.8 1.5 13.4 0.4 1.1
Burke Shaw 4.1 0.2 0.6 10.8 0.5 1.5 9.0 0.5 2.4 13.0 0.4 2.0 13.2 0.4 1.9 13.3 0.6 1.8 11.5 0.8 0.6 12.1 0.6 1.0 12.2 0.7 1.2 10.5 0.7 2.9 12.2 0.9 1.3 13.7 0.5 2.6
Ca Two Plus Quasip. 2.3 0.1 0.1 7.2 0.1 0.1 8.3 0.1 0.1 9.2 0.4 0.1 13.1 0.4 0.1 13.2 0.4 0.2 11.3 0.9 0.8 11.6 0.8 1.2 12.8 0.6 0.3 8.9 0.6 1.6 11.1 0.8 1.6 13.8 0.7 1.6
Chen Lee 3.9 0.4 0.3 4.9 0.3 0.9 4.0 0.5 1.6 4.8 0.3 0.4 4.6 0.3 1.0 8.9 0.7 1.3 7.2 0.9 3.6 7.9 0.7 2.1 8.0 0.7 1.7 4.8 0.5 1.4 11.0 0.9 1.2 6.4 0.5 4.0
Chua 5.9 0.1 0.1 7.4 0.2 0.2 5.9 0.2 0.2 7.6 0.1 0.2 11.7 0.2 0.1 13.1 0.3 0.4 11.8 0.6 1.0 12.2 0.5 1.3 11.7 0.6 0.6 12.0 0.6 1.4 11.3 0.5 1.5 13.4 0.4 0.8
Coullet 1.2 0.0 0.4 4.9 0.0 0.2 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.0 14.6 0.4 0.3 13.0 0.8 1.7 11.2 0.8 2.4 11.8 0.8 2.3 11.6 0.7 2.4 12.3 0.8 1.7 14.9 0.5 1.9
Dadras 4.1 0.2 1.5 7.7 0.6 0.9 4.7 0.5 1.2 13.1 0.5 0.5 12.1 0.4 0.3 10.2 0.7 1.0 12.7 0.8 2.6 12.9 0.7 2.4 13.0 0.8 2.7 8.8 0.8 4.2 11.9 0.8 2.5 12.8 0.4 2.8
Dequan Li 12.7 0.5 0.3 11.1 0.3 0.2 11.0 0.4 0.4 12.1 0.3 0.3 12.1 0.2 0.5 13.0 0.3 1.5 11.7 0.4 1.1 11.3 0.4 1.0 9.6 0.4 0.7 11.0 0.4 0.4 10.7 0.4 0.6 13.5 0.4 2.5
Finance 3.2 0.1 0.4 8.5 0.6 0.9 8.6 0.4 0.7 9.1 0.4 0.3 12.5 0.3 0.1 13.8 0.4 1.4 11.6 0.9 1.8 10.7 0.8 2.0 13.4 0.8 1.5 11.8 0.8 1.1 11.1 0.9 0.8 13.7 0.5 1.3
Genesio Tesi 8.2 0.2 0.3 7.9 0.4 0.7 8.5 0.2 0.7 7.7 0.2 0.7 9.4 0.2 0.8 13.0 0.3 1.6 10.4 0.7 0.7 10.6 0.7 0.8 10.3 0.8 1.0 9.6 0.8 1.5 11.0 0.9 0.9 12.6 0.6 4.1
Guckenheimer Holmes 9.6 0.1 0.2 11.8 0.1 0.4 12.0 0.1 0.4 12.6 0.3 0.2 12.5 0.4 0.6 12.1 0.4 0.2 9.6 0.7 0.2 9.6 0.7 0.4 9.4 0.6 0.2 11.8 0.6 0.4 10.2 0.7 0.5 12.7 0.5 0.6
Hadley 11.3 0.4 0.6 10.7 0.5 2.3 9.7 0.5 1.3 12.5 0.3 1.1 12.6 0.6 2.1 12.6 0.4 3.2 11.4 0.8 3.2 11.9 0.7 3.3 11.9 0.6 2.9 11.5 0.8 3.4 11.4 0.8 3.5 12.5 0.5 3.6
Halvorsen 4.3 0.3 1.0 9.7 0.6 0.3 8.2 0.2 0.8 7.5 0.2 0.8 13.1 0.2 0.6 13.3 0.4 1.3 10.0 0.8 4.7 10.3 0.7 4.4 9.7 0.8 4.4 10.9 0.7 4.3 11.3 0.9 3.4 13.4 0.5 3.5
Hindmarsh Rose 6.5 0.4 0.5 9.8 0.2 0.1 4.9 0.1 0.0 9.6 0.2 0.0 12.7 0.6 0.2 13.7 0.6 0.5 12.9 0.7 1.1 14.3 0.8 1.3 12.9 0.8 1.1 14.0 0.8 1.4 13.2 0.8 0.6 14.6 0.4 1.1
Isothermal Chemical 2.0 0.2 0.3 2.5 0.1 0.2 5.4 0.1 0.1 8.3 0.2 0.1 7.0 0.4 0.1 11.6 0.5 0.2 11.7 0.9 0.6 11.7 0.9 1.0 10.8 0.8 0.8 12.5 0.9 0.8 10.1 0.9 0.9 15.6 0.6 2.5
Itik Banks tumor 5.7 0.2 0.1 8.2 0.5 0.1 10.8 0.6 0.2 9.8 0.5 0.1 11.8 0.6 0.2 13.3 0.6 0.1 10.8 0.6 0.9 11.2 0.8 1.7 11.9 0.9 0.9 12.0 0.8 2.4 12.0 0.9 1.4 12.9 0.7 1.5
Jerk Circuit 4.3 0.3 0.4 8.8 0.6 1.1 9.1 0.4 0.8 10.6 0.3 1.0 11.7 0.3 0.2 9.8 0.6 1.1 12.5 0.9 1.9 13.5 0.9 1.6 12.9 0.9 0.7 13.1 1.0 1.5 10.3 0.8 0.5 14.1 0.5 3.1
Laser 4.5 0.2 0.5 7.5 0.5 5.0 8.2 0.3 1.9 11.5 0.2 0.6 12.8 0.2 0.7 13.9 0.6 1.3 12.6 0.7 2.4 12.8 0.5 1.7 13.0 0.7 1.3 12.8 0.7 2.7 11.8 0.8 2.4 13.3 0.5 2.2
Lorenz 3.4 0.2 0.6 9.8 0.6 0.5 8.5 0.3 0.6 13.0 0.4 0.7 12.6 0.7 0.8 13.3 0.6 0.3 12.8 0.8 1.0 13.0 0.7 1.9 13.2 0.6 1.4 13.0 0.6 1.3 11.8 0.8 1.3 13.4 0.5 1.1
Lorenz 84 7.9 0.2 1.6 10.7 0.5 3.3 11.0 0.4 4.4 12.6 0.4 4.8 11.0 0.4 0.8 13.2 0.6 1.8 11.8 0.9 2.0 12.2 0.9 2.6 11.5 0.9 2.4 12.3 0.9 2.5 11.2 0.9 2.2 13.4 0.6 2.9
Lu Chen 9.5 0.8 1.4 5.3 0.2 1.0 9.0 0.3 1.3 8.9 0.2 0.2 9.7 0.3 0.8 8.3 0.3 1.0 12.2 1.0 1.0 13.8 0.9 1.4 12.2 0.9 1.7 13.1 1.0 0.9 10.1 0.8 1.4 14.2 0.5 2.2
Moore Spiegel 9.2 0.6 0.9 9.1 0.6 0.9 10.9 0.7 0.9 12.5 0.5 0.8 11.5 0.6 0.8 11.3 0.4 0.9 11.4 0.9 2.6 11.7 0.8 3.6 12.7 0.9 3.6 12.0 0.8 3.3 10.5 0.9 4.6 13.1 0.7 4.0
Multi Chua 12.5 0.5 0.2 12.0 0.6 0.5 11.9 0.6 0.3 13.4 0.4 0.5 13.0 0.4 0.3 13.9 0.4 0.4 12.5 0.5 0.5 12.6 0.5 0.4 13.6 0.5 0.3 13.0 0.5 0.5 10.7 0.5 0.1 13.5 0.5 0.8
Nose Hoover 11.4 0.5 0.5 10.2 0.6 0.4 10.8 0.7 0.6 10.8 0.6 0.3 12.9 0.8 0.6 13.3 0.5 0.6 11.2 0.9 1.6 11.3 0.9 1.4 12.7 0.7 2.2 11.0 0.9 2.0 10.5 0.9 2.3 13.3 0.5 3.1
Pehlivan Wei 4.8 0.2 3.1 4.6 0.1 0.4 6.8 0.4 0.5 9.0 0.3 0.6 10.7 0.3 0.9 12.4 0.5 1.5 11.1 0.9 5.8 12.0 0.6 6.1 12.6 0.8 6.1 12.3 0.8 5.5 10.0 0.9 2.3 12.7 0.4 6.3
Sakarya 3.1 0.3 1.0 2.6 0.2 0.5 5.0 0.3 0.5 8.2 0.3 0.5 10.2 0.7 1.0 12.9 0.8 1.8 10.7 0.9 1.5 12.7 0.9 1.9 11.8 0.8 2.0 12.8 0.8 2.1 12.4 0.7 1.4 14.0 0.5 2.9
San Um Srisuchinwong 5.5 0.1 0.5 5.5 0.2 1.5 10.5 0.8 1.5 12.6 0.4 1.5 13.0 0.5 1.4 13.1 0.6 0.7 11.8 0.8 1.4 13.0 0.7 0.3 12.0 0.6 1.1 11.9 0.7 0.6 10.7 0.5 1.7 12.6 0.5 0.5
Sprott A 4.9 0.1 0.6 9.3 0.5 0.2 8.5 0.4 0.3 7.5 0.3 0.3 11.7 0.3 0.6 12.6 0.5 0.8 10.7 0.9 2.9 11.8 0.7 2.9 11.5 0.8 3.1 11.1 0.7 3.5 9.8 0.9 0.6 12.4 0.6 3.5
Sprott B 2.6 0.2 2.1 9.2 0.5 1.1 11.0 0.6 1.3 13.2 0.5 2.0 13.4 0.6 0.9 13.4 0.7 2.3 12.6 0.8 1.2 13.2 0.8 1.9 12.5 0.8 2.0 12.6 0.7 1.8 13.9 0.8 1.9 12.9 0.6 1.5
Sprott C 2.9 0.2 0.3 8.5 0.6 1.2 12.0 0.3 1.2 13.3 0.4 0.5 12.6 0.4 0.6 11.8 0.5 2.4 12.4 0.7 2.2 13.0 0.6 2.1 11.9 0.6 2.9 11.8 0.6 2.1 13.3 0.7 2.2 13.5 0.5 2.9
Sprott D 3.8 0.1 0.3 7.6 0.2 0.1 5.5 0.1 0.1 5.5 0.2 0.1 12.9 0.4 0.5 13.1 0.5 0.6 12.1 0.7 0.5 12.0 0.8 3.1 12.1 0.7 1.8 12.0 0.6 2.0 12.9 0.7 0.7 13.8 0.5 1.1
Sprott F 4.9 0.3 1.1 10.8 0.4 0.9 8.9 0.5 0.7 12.6 0.3 0.5 13.0 0.5 0.5 13.1 0.4 0.4 12.3 0.9 1.8 12.2 0.6 2.1 13.2 0.8 2.0 12.4 0.7 2.4 11.8 0.9 1.0 13.4 0.6 1.3
Sprott G 2.4 0.0 0.2 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.7 12.9 0.3 0.6 12.3 0.8 3.5 12.4 0.7 3.9 12.9 0.8 2.9 11.2 0.9 4.4 11.0 0.9 2.3 14.0 0.5 1.6
Sprott I 3.0 0.2 0.6 8.6 0.5 0.1 10.8 0.1 0.1 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.1 11.9 0.8 2.9 9.4 0.9 1.0 13.4 0.9 2.2 10.9 0.9 1.0 14.4 0.6 2.7
Sprott L 2.7 0.2 1.4 9.2 0.3 0.2 8.2 0.1 0.3 10.2 0.2 0.2 11.3 0.2 0.3 13.6 0.4 0.2 12.3 0.9 1.0 12.3 0.9 0.4 13.3 0.9 1.6 13.1 0.9 2.0 12.2 0.8 0.5 14.2 0.5 3.1
Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.1 12.8 0.3 0.3 13.5 0.3 0.3 14.0 0.8 0.5 12.1 0.8 0.9 11.1 0.7 1.3 13.0 0.8 0.9 12.3 0.8 0.8 10.7 0.9 0.8 14.3 0.6 2.3
Sprott O 3.5 0.2 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.6 11.6 0.4 0.4 13.2 0.3 0.5 11.8 1.0 0.9 12.7 0.8 1.6 11.4 1.0 1.8 12.6 1.0 3.9 10.4 0.9 0.9 14.0 0.7 3.5
Sprott P 5.7 0.2 1.2 9.2 0.6 0.8 10.5 0.4 0.5 11.1 0.8 1.0 13.3 0.4 0.7 13.5 0.6 1.6 9.6 0.8 4.0 7.6 0.6 3.3 8.7 0.8 3.4 9.6 0.8 3.0 10.8 0.8 1.5 10.4 0.5 1.8
Sprott Q 5.8 0.2 1.1 9.5 0.5 0.5 9.4 0.5 0.4 10.0 0.7 0.5 12.9 0.5 1.3 12.6 0.3 1.3 11.1 0.9 0.8 13.0 0.7 1.8 12.5 0.7 0.6 11.5 0.8 1.7 9.7 0.9 0.4 13.2 0.4 4.0
Sprott R 5.1 0.3 0.5 6.9 0.4 0.1 6.7 0.3 0.1 6.5 0.2 0.0 12.5 0.3 0.3 14.2 0.5 0.4 11.0 0.9 1.4 8.8 0.7 1.0 10.6 0.9 1.0 9.4 0.9 0.9 11.4 0.9 1.7 13.8 0.6 2.3
Sprott S 2.9 0.0 0.3 9.6 0.4 0.3 10.7 0.2 0.2 11.7 0.3 0.3 13.0 0.4 0.3 13.3 0.4 0.7 10.2 0.9 0.6 10.2 0.8 1.6 11.2 0.9 1.7 11.0 0.8 2.0 10.7 0.9 1.9 12.8 0.5 4.0
Sprott Torus 3.6 0.2 0.2 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.7 7.0 0.9 0.6 12.1 0.4 0.8 7.9 0.9 0.6 12.5 0.7 0.1 12.2 0.9 0.3 14.9 0.5 1.0
[…]

### v2

@@ A.7 Comparison of DynaMix to custom-trained DSR models
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 8.4 0.6 2.3 9.7 0.5 0.9 11.1 0.4 0.5 10.5 0.6 0.7 10.1 0.5 1.0 12.0 0.4 0.5 11.9 0.9 2.2 12.0 0.9 1.9 13.5 0.9 1.8 12.7 0.8 2.7 11.3 0.9 1.3 14.2 0.6 3.1 13.9 0.4 1.2
BeerRNN 11.7 0.8 1.2 8.0 0.5 0.2 3.3 0.2 0.2 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.2 12.5 0.9 1.2 12.1 0.9 1.2 12.6 0.8 1.1 13.0 0.9 1.4 14.1 0.6 1.3 14.0 0.4 0.8
Bouali 4.1 0.1 0.5 8.3 0.5 0.1 7.3 0.3 0.2 8.6 0.1 0.2 11.1 0.3 0.5 13.2 0.5 0.3 10.6 0.8 1.0 9.6 0.7 1.1 12.2 0.7 2.0 12.3 0.8 1.2 12.9 0.8 2.3 13.4 0.4 2.3 12.7 0.3 0.7
Burke Shaw 4.1 0.2 0.5 10.8 0.5 1.2 9.0 0.5 1.5 13.0 0.4 1.1 13.2 0.4 1.5 13.3 0.6 0.9 11.5 0.8 1.3 12.1 0.6 2.4 12.2 0.7 2.2 10.5 0.7 4.8 12.2 0.9 3.7 13.7 0.5 3.6 13.9 0.4 1.5
Ca Two Plus Quasip. 2.3 0.1 0.7 7.2 0.1 0.2 8.3 0.1 0.3 9.2 0.4 0.3 13.1 0.4 0.2 13.2 0.4 0.6 11.3 0.9 1.0 11.6 0.8 1.1 12.8 0.6 1.2 8.9 0.6 1.6 11.1 0.8 3.0 13.8 0.7 2.6 13.4 0.3 1.9
Chen Lee 3.9 0.4 0.3 4.9 0.3 0.3 4.0 0.5 0.5 4.8 0.3 0.1 4.6 0.3 0.3 8.9 0.7 0.4 7.2 0.9 1.6 7.9 0.7 1.4 8.0 0.7 2.0 4.8 0.5 1.7 11.0 0.9 0.9 6.4 0.5 3.4 7.3 0.4 1.8
Chua 5.9 0.1 0.3 7.4 0.2 0.6 5.9 0.2 0.7 7.6 0.1 0.7 11.7 0.2 0.4 13.1 0.3 1.4 11.8 0.6 1.9 12.2 0.5 2.2 11.7 0.6 1.7 12.0 0.6 2.7 11.3 0.5 3.2 13.4 0.4 2.7 10.4 0.2 0.7
Coullet 1.2 0.0 0.2 4.9 0.0 0.3 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.1 14.6 0.4 0.6 13.0 0.8 2.8 11.2 0.8 3.8 11.8 0.8 3.4 11.6 0.7 4.9 12.3 0.8 2.4 14.9 0.5 3.3 10.7 0.2 0.5
Dadras 4.1 0.2 1.1 7.7 0.6 0.5 4.7 0.5 0.5 13.1 0.5 0.3 12.1 0.4 0.2 10.2 0.7 0.5 12.7 0.8 1.7 12.9 0.7 2.0 13.0 0.8 2.2 8.8 0.8 3.8 11.9 0.8 1.7 12.8 0.4 2.9 11.8 0.4 1.2
Dequan Li 12.7 0.5 2.1 11.1 0.3 1.1 11.0 0.4 0.7 12.1 0.3 0.9 12.1 0.2 1.9 13.0 0.3 3.1 11.7 0.4 3.1 11.3 0.4 2.8 9.6 0.4 2.4 11.0 0.4 1.8 10.7 0.4 3.1 13.5 0.4 4.5 12.8 0.3 2.0
Finance 3.2 0.1 0.6 8.5 0.6 1.6 8.6 0.4 0.9 9.1 0.4 0.4 12.5 0.3 0.4 13.8 0.4 1.5 11.6 0.9 1.8 10.7 0.8 1.7 13.4 0.8 2.0 11.8 0.8 1.7 11.1 0.9 2.3 13.7 0.5 4.2 13.6 0.4 2.6
Genesio Tesi 8.2 0.2 0.9 7.9 0.4 1.0 8.5 0.2 1.1 7.7 0.2 1.0 9.4 0.2 1.0 13.0 0.3 2.0 10.4 0.7 1.4 10.6 0.7 1.1 10.3 0.8 2.2 9.6 0.8 2.1 11.0 0.9 1.4 12.6 0.6 5.4 10.6 0.4 1.0
Guckenheimer Holmes 9.6 0.1 0.5 11.8 0.1 0.8 12.0 0.1 0.4 12.6 0.3 0.5 12.5 0.4 1.1 12.1 0.4 0.6 9.6 0.7 0.6 9.6 0.7 2.5 9.4 0.6 1.0 11.8 0.6 1.4 10.2 0.7 1.0 12.7 0.5 3.5 11.5 0.3 0.9
Hadley 11.3 0.4 1.3 10.7 0.5 2.3 9.7 0.5 1.8 12.5 0.3 1.5 12.6 0.6 1.8 12.6 0.4 2.8 11.4 0.8 5.1 11.9 0.7 4.7 11.9 0.6 3.9 11.5 0.8 5.7 11.4 0.8 6.3 12.5 0.5 7.1 12.4 0.5 3.8
Halvorsen 4.3 0.3 0.6 9.7 0.6 0.4 8.2 0.2 0.7 7.5 0.2 0.7 13.1 0.2 0.6 13.3 0.4 1.9 10.0 0.8 5.0 10.3 0.7 4.3 9.7 0.8 4.6 10.9 0.7 4.1 11.3 0.9 3.4 13.4 0.5 4.2 13.3 0.3 1.6
Hindmarsh Rose 6.5 0.4 0.7 9.8 0.2 0.2 4.9 0.1 0.1 9.6 0.2 0.1 12.7 0.6 0.3 13.7 0.6 0.9 12.9 0.7 1.6 14.3 0.8 2.0 12.9 0.8 1.7 14.0 0.8 1.9 13.2 0.8 2.5 14.6 0.4 1.4 14.3 0.5 1.1
Isothermal Chemical 2.0 0.2 0.2 2.5 0.1 0.6 5.4 0.1 0.3 8.3 0.2 0.2 7.0 0.4 0.2 11.6 0.5 0.5 11.7 0.9 1.1 11.7 0.9 1.5 10.8 0.8 1.6 12.5 0.9 2.0 10.1 0.9 0.7 15.6 0.6 3.5 7.5 0.3 1.3
Itik Banks tumor 5.7 0.2 0.5 8.2 0.5 0.5 10.8 0.6 0.4 9.8 0.5 0.3 11.8 0.6 0.4 13.3 0.6 0.3 10.8 0.6 3.4 11.2 0.8 3.8 11.9 0.9 2.9 12.0 0.8 5.0 12.0 0.9 2.9 12.9 0.7 4.1 13.4 0.3 1.1
Jerk Circuit 4.3 0.3 0.8 8.8 0.6 0.4 9.1 0.4 0.3 10.6 0.3 0.2 11.7 0.3 0.2 9.8 0.6 0.3 12.5 0.9 1.0 13.5 0.9 0.8 12.9 0.9 0.8 13.1 1.0 0.9 10.3 0.8 0.6 14.1 0.5 1.9 13.7 0.4 1.1
Laser 4.5 0.2 0.2 7.5 0.5 3.0 8.2 0.3 1.2 11.5 0.2 0.5 12.8 0.2 0.5 13.9 0.6 0.8 12.6 0.7 2.0 12.8 0.5 2.8 13.0 0.7 1.7 12.8 0.7 5.1 11.8 0.8 2.5 13.3 0.5 2.6 13.8 0.4 1.0
Lorenz 3.4 0.2 0.5 9.8 0.6 1.5 8.5 0.3 1.6 13.0 0.4 0.7 12.6 0.7 2.0 13.3 0.6 0.8 12.8 0.8 2.6 13.0 0.7 3.4 13.2 0.6 3.0 13.0 0.6 2.3 11.8 0.8 2.7 13.4 0.5 3.1 13.1 0.3 2.1
Lorenz 84 7.9 0.2 0.5 10.7 0.5 1.5 11.0 0.4 1.3 12.6 0.4 1.0 11.0 0.4 0.4 13.2 0.6 0.4 11.8 0.9 1.3 12.2 0.9 1.4 11.5 0.9 1.3 12.3 0.9 1.3 11.2 0.9 1.2 13.4 0.6 1.2 13.4 0.5 1.4
Lu Chen 9.5 0.8 0.7 5.3 0.2 0.5 9.0 0.3 0.7 8.9 0.2 0.2 9.7 0.3 0.5 8.3 0.3 0.6 12.2 1.0 1.7 13.8 0.9 1.8 12.2 0.9 1.6 13.1 1.0 2.1 10.1 0.8 1.1 14.2 0.5 1.9 7.4 0.5 0.9
Moore Spiegel 9.2 0.6 0.4 9.1 0.6 0.4 10.9 0.7 0.4 12.5 0.5 0.3 11.5 0.6 0.3 11.3 0.4 0.3 11.4 0.9 2.0 11.7 0.8 2.7 12.7 0.9 2.0 12.0 0.8 2.4 10.5 0.9 2.2 13.1 0.7 1.5 12.6 0.5 2.1
Multi Chua 12.5 0.5 0.4 12.0 0.6 0.8 11.9 0.6 0.4 13.4 0.4 1.0 13.0 0.4 0.8 13.9 0.4 1.3 12.5 0.5 1.5 12.6 0.5 1.1 13.6 0.5 1.1 13.0 0.5 1.2 10.7 0.5 0.7 13.5 0.5 2.1 12.7 0.3 1.9
Nose Hoover 11.4 0.5 0.9 10.2 0.6 0.5 10.8 0.7 0.5 10.8 0.6 0.5 12.9 0.8 0.6 13.3 0.5 1.0 11.2 0.9 2.0 11.3 0.9 1.4 12.7 0.7 2.0 11.0 0.9 2.2 10.5 0.9 2.9 13.3 0.5 3.3 12.8 0.4 0.5
Pehlivan Wei 4.8 0.2 1.9 4.6 0.1 0.7 6.8 0.4 0.9 9.0 0.3 1.0 10.7 0.3 1.3 12.4 0.5 2.1 11.1 0.9 5.6 12.0 0.6 6.8 12.6 0.8 6.5 12.3 0.8 6.3 10.0 0.9 4.1 12.7 0.4 7.7 9.1 0.3 3.3
Sakarya 3.1 0.3 0.4 2.6 0.2 0.2 5.0 0.3 0.3 8.2 0.3 0.3 10.2 0.7 0.6 12.9 0.8 0.9 10.7 0.9 0.5 12.7 0.9 2.5 11.8 0.8 0.9 12.8 0.8 1.7 12.4 0.7 0.8 14.0 0.5 2.5 13.8 0.5 1.5
San Um Srisuchinwong 5.5 0.1 0.6 5.5 0.2 1.2 10.5 0.8 1.3 12.6 0.4 1.2 13.0 0.5 0.8 13.1 0.6 0.6 11.8 0.8 1.1 13.0 0.7 0.5 12.0 0.6 1.2 11.9 0.7 0.9 10.7 0.5 1.8 12.6 0.5 0.9 12.6 0.4 1.4
Sprott A 4.9 0.1 1.0 9.3 0.5 0.8 8.5 0.4 0.8 7.5 0.3 0.4 11.7 0.3 0.9 12.6 0.5 2.2 10.7 0.9 4.1 11.8 0.7 4.5 11.5 0.8 4.4 11.1 0.7 5.0 9.8 0.9 2.1 12.4 0.6 6.0 11.2 0.2 1.1
Sprott B 2.6 0.2 0.9 9.2 0.5 1.7 11.0 0.6 1.8 13.2 0.5 2.5 13.4 0.6 1.5 13.4 0.7 1.9 12.6 0.8 2.7 13.2 0.8 2.7 12.5 0.8 2.2 12.6 0.7 1.9 13.9 0.8 3.9 12.9 0.6 2.9 13.7 0.3 2.5
Sprott C 2.9 0.2 0.4 8.5 0.6 1.1 12.0 0.3 0.5 13.3 0.4 1.1 12.6 0.4 0.8 11.8 0.5 2.0 12.4 0.7 2.0 13.0 0.6 2.2 11.9 0.6 2.5 11.8 0.6 1.7 13.3 0.7 2.6 13.5 0.5 2.9 13.5 0.4 1.9
Sprott D 3.8 0.1 0.6 7.6 0.2 0.2 5.5 0.1 0.2 5.5 0.2 0.1 12.9 0.4 0.7 13.1 0.5 0.5 12.1 0.7 0.8 12.0 0.8 4.6 12.1 0.7 3.1 12.0 0.6 3.7 12.9 0.7 1.9 13.8 0.5 2.3 13.2 0.5 0.6
Sprott F 4.9 0.3 0.5 10.8 0.4 0.3 8.9 0.5 0.5 12.6 0.3 0.3 13.0 0.5 0.2 13.1 0.4 0.3 12.3 0.9 1.2 12.2 0.6 1.3 13.2 0.8 1.2 12.4 0.7 1.3 11.8 0.9 1.3 13.4 0.6 1.8 13.2 0.5 1.2
Sprott G 2.4 0.0 0.3 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.5 12.9 0.3 0.6 12.3 0.8 3.6 12.4 0.7 4.0 12.9 0.8 3.4 11.2 0.9 4.5 11.0 0.9 2.6 14.0 0.5 2.0 10.6 0.3 1.2
Sprott I 3.0 0.2 0.3 8.6 0.5 0.1 10.8 0.1 0.2 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.2 11.9 0.8 2.5 9.4 0.9 1.0 13.4 0.9 2.4 10.9 0.9 0.9 14.4 0.6 3.2 9.1 0.4 0.3
Sprott L 2.7 0.2 0.2 9.2 0.3 0.2 8.2 0.1 0.2 10.2 0.2 0.1 11.3 0.2 0.1 13.6 0.4 0.2 12.3 0.9 2.6 12.3 0.9 2.7 13.3 0.9 1.8 13.1 0.9 4.8 12.2 0.8 0.6 14.2 0.5 2.5 11.7 0.3 0.2
Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.2 12.8 0.3 0.3 13.5 0.3 0.6 14.0 0.8 0.7 12.1 0.8 1.9 11.1 0.7 2.3 13.0 0.8 1.2 12.3 0.8 1.5 10.7 0.9 1.3 14.3 0.6 3.3 13.5 0.4 0.6
Sprott O 3.5 0.1 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.4 11.6 0.4 0.3 13.2 0.3 0.5 11.8 1.0 1.2 12.7 0.8 2.1 11.4 1.0 2.3 12.6 1.0 3.1 10.4 0.9 0.8 14.0 0.7 3.2 9.6 0.2 0.5
Sprott P 5.7 0.2 0.4 9.2 0.6 0.9 10.5 0.4 0.6 11.1 0.8 0.9 13.3 0.4 0.5 13.5 0.6 1.7 9.6 0.8 6.8 7.6 0.6 5.7 8.7 0.8 4.7 9.6 0.8 5.8 10.8 0.8 1.4 10.4 0.5 3.2 9.1 0.5 1.9
Sprott Q 5.8 0.2 0.8 9.5 0.5 0.5 9.4 0.5 0.5 10.0 0.7 0.4 12.9 0.5 1.3 12.6 0.3 1.5 11.1 0.9 1.7 13.0 0.7 2.2 12.5 0.7 0.7 11.5 0.8 3.8 9.7 0.9 1.1 13.2 0.4 6.0 12.3 0.4 0.8
Sprott R 5.1 0.3 0.7 6.9 0.4 0.3 6.7 0.3 0.2 6.5 0.2 0.1 12.5 0.3 0.7 14.2 0.5 1.4 11.0 0.9 2.8 8.8 0.7 2.2 10.6 0.9 3.2 9.4 0.9 1.9 11.4 0.9 2.2 13.8 0.6 6.0 8.9 0.3 1.6
Sprott S 2.9 0.0 0.3 9.6 0.4 0.4 10.7 0.2 0.3 11.7 0.3 0.5 13.0 0.4 0.2 13.3 0.4 0.6 10.2 0.9 0.8 10.2 0.8 3.4 11.2 0.9 3.1 11.0 0.8 3.5 10.7 0.9 2.2 12.8 0.5 6.0 11.3 0.4 1.9
Sprott Torus 3.6 0.2 0.1 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.8 7.0 0.9 1.1 12.1 0.4 1.1 7.9 0.9 0.8 12.5 0.7 0.8 12.2 0.9 1.1 14.9 0.5 1.5 12.8 0.4 0.9
[…]